# Distributed MIMO Networks with Rotary ULAs for Indoor Scenarios under Rician Fading

### v1

@@ IV-A Particle Swarm Optimization
PSO [40, Ch. 16] is very suitable for solving problems where a function’s global maximum or minimum is very difficult to find. The algorithm works over a population of candidate solutions, known as agents or particles, and moves them in the search space according to their current positions and velocities. Each particle’s movement is influenced by its best-known position and the global best-known position in the search space. The local and global best positions are updated on each iteration, and the algorithm is expected to move the swarm of particles toward the optimal solution.
@@
The parameters of the PSO algorithm are listed in Table V, while its pseudo-code is in Algorithm 1. The inertial weight $w$ controls the particle’s tendency to continue in its current direction, while $c_{1}$ and $c_{2}$ are the acceleration coefficients that control the influence of the personal and global best positions, respectively. In our optimization problem, the number of variables to be jointly optimized equals the number of APs, $Q$. Each particle is a candidate set of angular positions for the $Q$ RULAs, i.e. $\{\theta_{1},\ldots,\theta_{Q}\}$, which corresponds to a point in the $Q$-dimensional space with boundaries $[0,\pi]^{Q\times 1}$. The objective function depends on the estimates of the positions of the MTDs, i.e., $f(\hat{\textbf{p}}_{1},\ldots,\hat{\textbf{p}}_{K})$, and is described in the next Subsection. There are two possible termination criteria: i) the maximum number of iterations is reached, or ii) the relative change in the value of the global best over a predefined number of maximum stall iterations is less than the tolerance $\epsilon$.
TABLE II: Parameters of the PSO Algorithm
@@ IV-B Location-Based Beamforming
Before Phase 3 of the three-phase RA scheme, the CPU does not have any CSI information. Thus, the only information available for the computation of the optimal angular positions of the RULAs is the estimates of the positions of the MTDs that will be active in each time slot. Thus, we adopt a location-based beamforming [47, 48, 49, 50] approach to estimate the corresponding objective function value.
@@
Figure 4: The four different setups of APs are considered in this work.
@@ V Numerical Results
In this section, we present Monte Carlo simulation results that evince the performance improvements obtained by APs equipped with the proposed RULAs when compared to APs equipped with static ULAs. We keep the total number of antenna elements constant and study the trade-off between the number of APs and antenna elements per AP. The four different setups considered for the simulations are illustrated in Fig. 4.
@@ V-A Simulation Parameters
The F-quantile of a Rayleigh distribution with scaling parameter $\sigma_{e}$ is given by
$Q(F;\sigma_{e})=\sigma_{e}\sqrt{-2\ln(1-F)}.$ (27)
Considering the proposed positioning error model and based on the 3GPP specifications, we list examples of positioning accuracy requirements for pessimistic, reasonable, and optimistic scenarios and their corresponding values of $\sigma_{e}^{2}$, in Table V.
TABLE III: Simulation parameters.
@@
TABLE IV: Parameters of the PSO Algorithm.
@@
TABLE V: Examples of positioning accuracy requirements for pessimistic, optimistic, and reasonable scenarios [35, 36, 37].
$\phantom{-}\sigma_{e}^{2}$ 95% Quantile Scenario
[…]
$-7.8$ dB $1$ m Reasonable
$-28$ dB $10$ cm Optimistic
$-45$ dB $1$ cm Optimistic
@@ V-B Simulation Results and Discussions
The mean per-user achievable SE is obtained by averaging over $32$ network realizations, i.e., distinct sets of positions for the $K$ active MTDs. For each network realization, the achievable SE of the $K$ MTDs is obtained by averaging over $250$ channel realizations, i.e., distinct realizations of the channel matrix H.
@@
The mean-per user achievable SE $\bar{R}$ as a function of the Rician factor $\kappa$, for different values of $Q$ and $l$, are shown in Fig. 6. We observe that, in the case of APs equipped with static ULAs, the mean per-user achievable SE decreases with the Rician factor for any values of $Q$ and $l$. This happens because the correlation among the channel vectors increases with $\kappa$, thus affecting the performance obtained with the ZF combiner. On the other hand, in the case of APs equipped with the proposed RULAs, the mean per-user achievable SE increases with $\kappa$, excepted for the setups with $Q=8$. This evinces the fact that the optimal rotation of the RULAs reduces the correlation between the channel vectors, consequently enhancing the performance obtained with ZF when the LoS component becomes strong. Note that, when $\kappa\leq-10$ dB, i.e., when the channel tends to be pure NLoS, the optimal rotation of the RULAs brings no performance improvements, as expected. Nevertheless, when $\kappa$ grows large, the performance gains obtained with the RULAs become very significant for any values of $Q$ and $l$. Moreover, when the channel tends to be pure NLoS, increasing $Q$ enhances the mean per-user achievable SE for any value of $l$. Thus, in rich scattering environments, the most distributed D-MIMO setup with $Q=8$ is always the best choice, and the setup with $Q=4$ follows closely behind. In contrast, when $\kappa$ grows large and the LoS component becomes strong, there exists an optimal number of APs (and of antenna elements per AP) that maximizes the mean per-user achievable SE. In the case of APs with static ULAs, the best performance is achieved by the setup with $Q=2$ in the case of $l=50$ m, and by the setups with $Q=4$ for the cases of $l=100$ m and $l=200$ m. When we adopt APs equipped with RULAs, the single AP setup achieves the best performance for the case of $l=50$ m. When $l=100$ m, the setups with $Q=[1,2,4]$ achieve very similar performance. Finally, when $l=200$ m, the best performance is achieved when $Q=4$.
@@
The numerical results in Fig. 6 show that, in scenarios where the LoS component is strong, there is a sweet spot between the beamforming gains obtained with APs equipped with multiple antennas, and the macro-diversity gains obtained with the spatial distribution of APs. In the case of APs equipped with static ULAs, the intermediate deployments with $Q=[2,4]$ achieve this sweet spot. Conversely, when the proposed RULAs are adopted, utilizing a single AP allow us to achieve a satisfactory performance, while at the same time avoiding the higher deployment and maintenance costs of D-MIMO networks.
In Fig. 6, we show the mean per-user achievable SE versus the variance of the positioning error $\sigma_{e}^{2}$, for different values of $Q$ and $l$, and considering $\kappa=10$ dB. Interestingly, we observe that the performance gains owing to the optimal rotations of the RULAs are still very significant even when the accuracy of the position estimates is poor. Nevertheless, the SE is almost constant for $\sigma_{e}^{2}\leq-10$ dB, that is, improving the accuracy of the position estimates does not yield noticeable performance improvements.
Figure 5: Mean per-user achievable SE versus Rician factor, considering centralized ZF and $\sigma_{e}^{2}=-10$ dB, for (a) $l=50$ m, (b) $l=100$ m and (c) $l=200$ m.
Figure 6: Mean per-user achievable SE versus the variance of the positioning error, considering centralized ZF and $\kappa=10$ dB, for (a) $l=50$ m, (b) $l=100$ m and (c) $l=200$ m.
@@ VI Conclusions
In this work, we studied the performance of D-MIMO networks for indoor scenarios. Aiming to enhance the quality of the wireless links, we proposed the deployment of APs equipped with RULAs. The CPU jointly computes the optimal rotation of the RULAs as a function of the estimated positions of the active MTDs. The optimization is done using a location-based beamforming approach and PSO. We considered a spatially correlated Rician fading model and evaluated the setups’ performance for different Rician factor values. We took into account the impact of imperfect positioning information and also of the imperfect CSI. Given the total number of antenna elements, we evaluated the trade-off between the number of APs and antennas per AP. The numerical results show that the optimal rotation of the RULAs can bring substantial performance gains in terms of mean per-user achievable SE, and the gains grow with the Rician factor. We also observed that the optimal rotation of the RULAs can bring performance gains even when the positioning accuracy is poor or moderate. We also concluded that, given the total number of antenna elements, there is a sweet spot between the number of APs and the number of antenna elements per AP, corresponding to a trade-off between beamforming and macro-diversity gains. The best performance is achieved by adopting a few APs equipped with multiple antennas. This approach also has the advantage of reducing the deployment and maintenance costs of the system and its computational complexity.

### v2

@@ IV-A Particle Swarm Optimization
PSO [49, Ch. 16] is very suitable for solving problems where a function’s global maximum or minimum is very difficult to find. The algorithm works over a population of candidate solutions, known as agents or particles, and moves them in the search space according to their current positions and velocities. Each particle’s movement is influenced by its best-known position and the global best-known position in the search space. The local and global best positions are updated on each iteration, and the algorithm is expected to move the swarm of particles toward the optimal solution.
@@
The parameters of the PSO algorithm are listed in Table III, while its pseudo-code is in Algorithm 1. The inertial weight $w$ controls the particle’s tendency to continue in its current direction, while $c_{1}$ and $c_{2}$ are the acceleration coefficients that control the influence of the personal and global best positions, respectively. In our optimization problem, the number of variables to be jointly optimized equals the number of APs, $Q$. Each particle is a candidate set of angular positions for the $Q$ RULAs, i.e. $\{\theta_{1},\ldots,\theta_{Q}\}$, which corresponds to a point in the $Q$-dimensional space with boundaries $[0,\pi]^{Q\times 1}$. The objective function depends on the estimates of the locations of the MTDs, i.e., $f(\hat{\textbf{p}}_{1},\ldots,\hat{\textbf{p}}_{K})$, and is described in the next Subsection. There are two possible termination criteria: i) the maximum number of iterations is reached, or ii) the relative change in the value of the global best over a predefined number of maximum stall iterations is less than the tolerance $\epsilon$.
TABLE III: Parameters of the PSO Algorithm
@@
PSO is suitable for the obtaining the maximum global or minimum of functions since its particles evaluate the function at multiple points at the same. Thus, the optimization problem is solved in a parallelized fashion, which significantly reduces the required computation time. Besides, having multiple particles distributed on the domain of the function makes it harder for the problem to converge to a local minimum or maximum point. Nevertheless, the major drawbacks of PSO are [56, 57, 58]:
  1. 1.
    There is no guarantee that it will indeed converge to the global minimum or maximum point because of its stochastic nature. The particles may become clustered around a local optimum point early in the optimization process, leading to premature convergence;
  2. 2.
    Depending on the problem, PSO may be very sensitive to the tuning of its parameters, such as the number of the particles, inertia weight, cognitive and social constants, and the maximum number of iterations;
  3. 3.
    PSO may present a very high computational cost when dealing with high-dimensions optimization problems (i.e. problems with many optimization variables) and not perform well at the same time.
@@ IV-B Location-Based Beamforming
Before Phase 3 of the three-phase RA scheme, the CPU does not have any CSI information. Thus, the only information available for the computation of the optimal angular positions of the RULAs is the estimates of the locations of the MTDs that will be active in each time slot. Thus, we adopt a location-based beamforming [59, 60, 61, 62] approach to estimate the corresponding objective function value.
@@
Figure 4: The square coverage area with $l=50$ m and the four different setups of APs considered in this work.
@@ V Numerical Results
In this section, we present Monte Carlo simulation results that evince the performance improvements obtained by APs equipped with the proposed RULAs when compared to APs equipped with static ULAs. We keep the total number of antenna elements constant and study the trade-off between the number of APs and antenna elements per AP. The square coverage area [^6] We opted to consider a square coverage area keeping in mind the shape of indoor industrial environments such as factory floors and warehouses. A square coverage area also facilitates the simulations, since it is simpler to uniformly distributed the APs this way. Moreover, we assuming circular or hexagonal areas should not lead to any relevant changes on the numerical results and conclusions. It should also be noted that in real-world scenarios, coverage areas can have irregular shapes. with $l=50$ m and the four different setups of APs considered for the simulations are illustrated in Fig. 4.
@@ V-A Simulation Parameters
TABLE IV: Simulation parameters.
@@
TABLE V: Parameters of the PSO Algorithm.
@@
(a) $l=50$ m
(b) $l=100$ m
(c) $l=200$ m
Figure 5: Mean per-user achievable SE versus Rician factor, considering centralized ZF and $\sigma_{e}^{2}=-10$ dB, for different sizes of the square coverage area.
(d) $l=50$ m
(e) $l=100$ m
(f) $l=200$ m
Figure 6: Mean per-user achievable SE versus the variance of the localization error, considering centralized ZF and $\kappa=10$ dB, for different sizes of the square coverage area.
@@ V-B Simulation Results and Discussions
The mean per-user achievable SE is obtained by averaging over multiple network realizations, i.e., distinct sets of locations for the $K$ active MTDs. For each network realization, the achievable SE of the $K$ MTDs is obtained by averaging over several channel realizations, i.e., distinct realizations of the channel matrix H.
@@
The mean-per user achievable SE $\bar{R}$ as a function of the Rician factor $\kappa$, for different values of $Q$ and $l$, are shown in Fig. 6. The primary motivation for obtaining $\bar{R}$ as a function of $\kappa$ is to evaluate the performance of our proposed approach under a variety of different conditions. Different indoor industrial scenarios may feature diverse objects (varying in number, size and materials), the presence or absence of walls, leading to wireless channels that experience different Rician factors. We observe that, in the case of APs equipped with static ULAs, the mean per-user achievable SE decreases with the Rician factor for any values of $Q$ and $l$. This happens because the correlation among the channel vectors increases with $\kappa$, thus affecting the performance obtained with the ZF combiner. On the other hand, in the case of APs equipped with the proposed RULAs, the mean per-user achievable SE increases with $\kappa$, excepted for the setups with $Q=8$. This evinces the fact that the optimal rotation of the RULAs reduces the correlation between the channel vectors, consequently enhancing the performance obtained with ZF when the LoS component becomes strong. Note that, when $\kappa\leq-10$ dB, i.e., when the channel tends to be pure NLoS, the optimal rotation of the RULAs brings no performance improvements, as expected. Nevertheless, when $\kappa$ grows large, the performance gains obtained with the RULAs become very significant for any values of $Q$ and $l$. Moreover, when the channel tends to be pure NLoS, increasing $Q$ enhances the mean per-user achievable SE for any value of $l$. Thus, in rich scattering environments, the most distributed D-MIMO setup with $Q=8$ is always the best choice, and the setup with $Q=4$ follows closely behind. In contrast, when $\kappa$ grows large and the LoS component becomes strong, there exists an optimal number of APs (and of antenna elements per AP) that maximizes the mean per-user achievable SE.
@@
The numerical results in Fig. 6 show that in the case of APs with static ULAs and under high Rician factor, the best performance is achieved by the setup with $Q=2$ in the case of $l=50$ m (Fig. 6a), and by the setups with $Q=4$ for the cases of $l=100$ m and $l=200$ m (Figs. 6b and 6c, respectively). Conversely, when we adopt APs equipped with RULAs in the scenarios with high Rician factor, the single AP setup achieves the best performance for the case of $l=50$ m (Fig. 6a). When $l=100$ m (Fig. 6b), the setups with $Q=[1,2,4]$ achieve nearly identical performance. Finally, when $l=200$ m (Fig. 6c), the best performance is achieved when $Q=4$. Overall, the curves in Fig. 6 show that, when the LoS component is strong, there is a sweet spot between the beamforming gains obtained with APs equipped with multiple antennas, and the macro-diversity gains obtained with the spatial distribution of APs. In the case of APs equipped with static ULAs, the intermediate deployments with $Q=[2,4]$ achieve this sweet spot. Conversely, when the proposed RULAs are adopted, utilizing a single AP allows us to achieve satisfactory performance while avoiding D-MIMO networks’ higher deployment and maintenance costs.
In Fig. 6, we show the mean per-user achievable SE versus the variance of the localization error $\sigma_{e}^{2}$, for different values of $Q$ and $l$, and considering $\kappa=10$ dB. Note that all the curves in Fig. 6 corresponding to the case of static APs are constant since those setups do not use the estimated locations of the active MTDs for any optimization. As discussed in Section III, the scenarios where $\sigma_{e}^{2}\geq 5$ dB, that is, with very poor location estimates, are not expected to occur in current or future wireless networks under normal operational conditions. We evaluate this range of variance of the localization error to illustrate the robustness of our proposed framework. Interestingly, we observe that the performance gains owing to the optimal rotations of the RULAs are still very significant even when the accuracy of the location estimates is poor. Nevertheless, the SE is almost constant for $\sigma_{e}^{2}\leq-10$ dB, that is, improving the accuracy of the location estimates does not yield noticeable performance improvements. The curves in Fig. 6a show that, in the case of a small coverage area (specifically, $l=50$ m), it is more advantageous to deploy fewer APs equipped with more antennas since the best performance is achieved with $Q=2$ in the case of static ULAs and with $Q=1$ in the case of the proposed RULAs. When we increase the size of the coverage area to $l=100$ m and $l=200$ m (Figs. 6b and 6c, respectively), we observe that the setup with $Q=4$ achieves the best performance for both cases of static ULAs and RULAs. In the case of $l=100$ m, the performance achieved with $Q=1$ and $Q=2$ when RULAs are adopted approach the performance obtained with $Q=4$. These numerical results show that, as we increase the size of the coverage area, and given a total number of antenna elements $M$, a sweet spot between the number of APs and the number of antenna elements per AP becomes evident. This sweet spot corresponds to a trade-off between macro-diversity and beamforming gains. Finally, it is also very interesting to observe that in all the cases ($l=\{50,100,200\}$ m), the most distributed setup with $Q=8$ presents the worst performance.
V-C Complexity Analysis of PSO
The simulations ran on a personal computer having Windows 10, Intel(R) Core(TM) i5-10210U processor and 32 GB of RAM memory. Besides, we utilized Matlab version R2022b. The PSO was implemented on Matlab using the particleswarm built-in function [63]. Since we utilized a built-in function of Matlab, it is very difficult to measure the complexity of the PSO in terms of metrics such as number of float-point operations (FLOPs). Thus, we decided to analyse the computational complexity of our PSO solution in terms of the mean and the variance of the required time for the optimization algorithm to finish. This time can be measured in Matlab using the functions tic;toc or cputime.
Number of RULAs $Q=1$ $Q=2$ $Q=4$ $Q=8$
Mean 0.2652 s 0.6042 s 2.1520 s 7.9548 s
Variance 0.0363 s2 0.0636 s2 0.4624 s2 7.3269 s2
TABLE VI: Mean and variance of the execution time required for PSO to finish, considering $\kappa=10$ dB, $\sigma_{e}^{2}=-10$ dB and $l=100$ m.
The mean and variance of the execution time can be found in Table VI. The PSO converges when the relative change in the value of the objective function over a given number of “maximum stall iterations” is less than the “function tolerance” [63]. We observe that the mean required time increases rapidly with the number of RULAs $Q$ (i.e., the number of optimization variables). Therefore, PSO indeed may not be the best choice in D-MIMO networks with multiple RULAs whose angular positions are jointly optimized by the CPU. Moreover, the mean required time for any given value of $Q$ indicates that it is impractical to optimize the angular positions of the RULAs at every time slot. In other words, the angular positions should not be adjusted too often. In the framework we propose, the RULAs’ angular positions are adjusted a single time, prior to the uplink data transmission phase. Subsequently, the angular positions remain unchanged throughout multiple time slots during the uplink data transmission phase.
@@ VI Conclusions
In this work, we studied the performance of D-MIMO networks for indoor scenarios. Aiming to enhance the quality of the wireless links, we proposed the deployment of APs equipped with RULAs. The CPU jointly computes the optimal rotation of the RULAs as a function of the estimated locations of the active MTDs. The optimization is done using a location-based beamforming approach and PSO. We considered a spatially correlated Rician fading model and evaluated the setups’ performance for different Rician factor values. We took into account the impact of imperfect localization information and also of the imperfect CSI. Given the total number of antenna elements, we evaluated the trade-off between the number of APs and antennas per AP. The numerical results show that the optimal rotation of the RULAs can bring substantial performance gains in terms of mean per-user achievable SE, and the gains grow with the Rician factor. We also observed that the optimal rotation of the RULAs can bring performance gains even when the localization accuracy is poor or moderate. We also concluded that, given the total number of antenna elements, there is a sweet spot between the number of APs and the number of antenna elements per AP, corresponding to a trade-off between beamforming and macro-diversity gains. The best performance is achieved by adopting a few APs equipped with multiple antennas. This approach also has the advantage of reducing the deployment and maintenance costs of the system and its computational complexity.
This work studied the performance gains in terms of SE that the rotation of the ULAs can bring to wireless communications. Nevertheless, Integrated Sensing and Communications (ISAC) is a research topic that has become extremely popular in both academia and industry. Some recent works [64, 65, 66] have evaluated the performance improvements in the accuracy of sensing and localization obtained with the movement of antennas. The movement of antennas can be beneficial in several ways. For instance, it can help overcome blockages in the line-of-sight between an access point (AP) and a device. Additionally, by allowing for precise beamforming and optimal beamwidth design, movable antennas can enhance data transmission and improve sensing accuracy. This capability reduces undesirable side lobes and improves interference management. Furthermore, movable antennas can dynamically switch between optimal array geometries, depending on whether they are used for sensing or communication applications. Comparing different implementations of movable and/or rotary antennas and the associated performance improvements for both sensing and communications is a promising direction for future works.